OpenStudy (anonymous):

What does it mean for a number to be the "solution" to an equation? Give an example of a variable equation and its solution. Explain in complete sentences how you know that number is the solution

OpenStudy (anonymous):

What does it mean for a number to be the "solution" to an equation? Give an example of a variable equation and its solution. Explain in complete sentences how you know that number is the solution

OpenStudy (anonymous):

Let's say you have an equation \[x + 2 = 3\] The equation basically tells you that a variable (x) + 2 = 3 A variable doesn't have any defined value, theoretically it can take on any value. However, in this case, the equation can only be true, if \[x +2 -2 = 3 -2\] \[x = 1\] So you know that the variable x must be 1, in order for the equation to be true.